OpenStudy (anonymous):

Evaluate the expression when r = 8.2 and s = –3.7. r – 2s A. –22.94 B. 0.8 C. 4.5 D. 15.6

OpenStudy (anonymous):

@Hotchellerae21 @mathmate @Learner11

OpenStudy (anonymous):

i keep on failing this ;-;

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

r=8.2 s=-3.7

OpenStudy (♪chibiterasu):

Jsut plug them in

OpenStudy (♪chibiterasu):

8.2 - 2(-3.7)

OpenStudy (anonymous):

r – 2s =8.2-2(-3.7) =8.2 +7.4 =15.6

OpenStudy (anonymous):

So, it's D

OpenStudy (carlyleukhardt):

you should change your profile picture. you show to much in it. if you don't i might report you

OpenStudy (anonymous):

Um i didnt know my mom was on this site xD no one here is talking about profile pictures @carlyleukhardt can u please get off my question??

OpenStudy (♪chibiterasu):

It's fine, she's not showing any private body parts or sex organs and her pose is not sexual in any way.

OpenStudy (carlyleukhardt):

Well okay then.... stop tryng to get attention from boys..

OpenStudy (♪chibiterasu):

It's a normal selfie. :s

OpenStudy (carlyleukhardt):

shes trying to be sexy though...

OpenStudy (♪chibiterasu):

Who isn't?

OpenStudy (carlyleukhardt):

im just giving you a warning.

OpenStudy (♪chibiterasu):

Her pic is fine.

OpenStudy (anonymous):

Ya its fine

OpenStudy (♪chibiterasu):

All I recommend is changing the school name.

OpenStudy (♪chibiterasu):

Before you get a warning for bypassing our censors~

OpenStudy (carlyleukhardt):

yeah me to

OpenStudy (anonymous):

Yes i will change that but can you get off my profile? It kinda sounds like your jealous idk

OpenStudy (anonymous):

@carlyleukhardt

OpenStudy (kendricklamar2014):

Since this is all resolved. How about we close this question and live our lives

OpenStudy (carlyleukhardt):

im not jealous, i love myself, i mean kinda...